The probability of rolling exactly three 4's in 5 rolls of a fair die
Found 3 solutions by ewatrrr, stanbon, edjones:
Answer by ewatrrr(24785)

 
Hi,
The probability of rolling exactly three 4's in 5 rolls of a fair die
P(rolling a four)= 1/6
P(not rolling a four) = 5/6
1/6*1/6*1/6*5/6*5/6
distinguishable orders of the five rolls/3fours and '2' others:5!/[3!*2!]=10
P(exactly 3 fours in five rolls) = 10*(1/6*1/6*1/6*5/6*5/6)

Answer by stanbon(75887)

The probability of rolling exactly three 4's in 5 rolls of a fair die
---
Binomial Problem with n = 5 and p = 1/6
---
P(x = 3) = 5C3(1/6)^3(5/6)^2 = 10*(25/6^5) = 0.0322
==================
Cheers,
Stan H.
=========

Answer by edjones(8007)

p=probability of rolling a four
p=1/6, q=5/6
(p+q)^5 expand
5C3=10 coefficient
10p^3q^2=.03215 third term of the expansion.
.
Ed